OpenStudy (anonymous):

sieben Sie Folge an , die .. 1) ... beschränkt sind , aber nicht konvergieren 2) ... konvergieren , aber nicht monoton steigend sind 3) ... Nulfolgen sind , deren Summen aber nicht konvergieren 4) ... weder nach oben noch nach unten beschränkt sind * I think this is the right translation but I am not sure :- Give an example of each of the following kinds of series : 1. bounded but divergent 2. convergent and not increasing 3. zero sequences, their sums diverge 4. are neither bounded above ( inceasing ), or below ( decreasing )

OpenStudy (anonymous):

OK so if I'm right, you used German on the top. Well let me help with your question in my next post...

OpenStudy (anonymous):

Let me see here... A bounded series that is not divergent? Maybe this one: (1, 0, 1, 0) Did this help? OK next up...

OpenStudy (anonymous):

I believe the correct translation was 'set', by the way.

OpenStudy (anonymous):

A convergent set can be any set that adds up to a limit. For example, you could say 1/2 + 1/2 + 1/2 + 1/2 = 4. I think you might be confused with the language I'm using, if not could you reply?

OpenStudy (anonymous):

Sorry, I do not speak English well. But I'm trying to understand you .. there is no problem if you continue in english language

OpenStudy (anonymous):

OK sorry my internet keeps going down. Now where was I...

OpenStudy (anonymous):

No problem :)

OpenStudy (anonymous):

-1/2 - - -1/2 - -1/2 - -1/2 = So ein gutes Beispiel für eine konvergente und einem, der nicht zusätzlich in es wäre, die gleiche Sache nur wie -1/2 sein 4.

OpenStudy (anonymous):

OK I am sincerely sorry about that I have a glitch translator so I am going to continue in English. Anyways, a convergent sequence that is not directly addition could me (-1/2 - -1/2 - -1/2 - -1/2 - -1/2) = 4

OpenStudy (anonymous):

OK a zero sequence whose sum diverges could be...

OpenStudy (anonymous):

I'm sorry but I couldn't find anything on that and I do not remember anything about divergent zero sequences, but I will help with the rest of your problem.

OpenStudy (anonymous):

OK this one should be fairly easy neither bounded nor below because there is a sequence known as pi whose digits are completely random and neither increase nor decrease constantly: (3, 1, 4, 1, 5, 9, 2, 6, 5)

OpenStudy (anonymous):

I hope this helps Rama94 and I hope you have a great time! :)

OpenStudy (anonymous):

Certainly, your answers helped me to understand the issue more .. Thank you very much :) I hope you have a great time too :)